E7-5 Calculating Cost of Ending Inventory and Cost of Goods Sold under Periodic FIFO, LIFO, and Weighted Average Cost [LO 7-3 just answer requird 3
Answer #1
Units Unit cost Total
Jan 1 120 80 9600
Jan 15 380 90 34200
Jan 24 200 110 22000
Total 700 65800
Average cost 94 =65800/700
Ending inventory units 460 =700-240
3
FIFO:
Cost of ending inventory 45400 =(200*110)+(460-200)*90
Cost of goods sold 20400 =65800-45400
LIFO:
Cost of ending inventory 40200 =(120*80)+(460-120)*90
Cost of goods sold 25600 =65800-40200
Weighted Average cost:
Cost of ending inventory 43240 =460*94
Cost of goods sold 22560 =65800-43240